OpenStudy (kylexchaos):

Medal and Fan! What is the LCD (Least Common Denominator) of 3/10 and -4/15?

OpenStudy (mathmale):

What are your two different denominators here? List them, please.

OpenStudy (kylexchaos):

I do not know, that's why I'm asking for help...

OpenStudy (kylexchaos):

so..

OpenStudy (mathmale):

You have two fractions. Every fraction has a denominator. I'm sure you can answer my question. Simply look at the 2 given fractions and identify the number on the bottom of each.

OpenStudy (kylexchaos):

10 and 15, right?

OpenStudy (mathmale):

yes. Now factor 10 as far as possible. What are the factors? Now factor 15 as far as poss. What are the factors of 15?

OpenStudy (kylexchaos):

10 = 1, 2, 5,

OpenStudy (kylexchaos):

right?

OpenStudy (mathmale):

yes.

OpenStudy (mathmale):

15=

OpenStudy (kylexchaos):

Then the factor of 15 is 1,3,5

OpenStudy (mathmale):

good. so you have 10=2*5 and 15=3*5. We see 5 as a factor of both. Begin writing your L:CD now. start with 5. What other factors of 10 and 15 do you see?

OpenStudy (kylexchaos):

I don't understand, sorry.

OpenStudy (mathmale):

The idea is to take all of the whole number factors of 10 and 15 and multiply them together. Both 10 and 15 have 5 as a factor, so the LCD contains 5. 10 also contains 2, so take it; multiply 5 by 2; similarly 15 contains 3, take it and multiply your 5*2 by 3. 5*2*3=?

OpenStudy (mathmale):

(5)(2)(3) = ?

OpenStudy (mathmale):

@KylexChaos ?

OpenStudy (mathmale):

I am trying to help you understand where LCDs come from. This is an essential skill. Do you want to continue, or not?

OpenStudy (kylexchaos):

30!

OpenStudy (kylexchaos):

@mathmale